OpenStudy (quickstudent):

Can someone please help me with this question? solutions to rational equations

OpenStudy (quickstudent):

When I was solving, the only solution I got was 5. But when I plug that into the equation, it comes out undefined. So I'm confused if 5 is an extraneous solution, or if the equation has no solution at all. Explain?

OpenStudy (mathstudent55):

Let's start by factoring the right denominator. \(\dfrac{1}{r - 5} = \dfrac{5}{5(25 - 5r)} \) \(\dfrac{1}{r - 5} = \dfrac{5}{5\times 5(5 - r)} \) \(\dfrac{1}{r - 5} = \dfrac{\cancel{5}~1}{\cancel{5\times }5(5 - r)} \) \(\dfrac{1}{r - 5} = \dfrac{1}{5(5 - r)} \) Look at the two denominators now to find restrictions in the domain. \(r - 5 = 0\) and \(5 - r = 0\) The restrictions are both the same, \(r \ne 5\)

OpenStudy (mathstudent55):

Now we can solve the equation using cross multiplication. \(\dfrac{1}{r - 5} = \dfrac{1}{5(5 - r)} \) \(r - 5 = 5(5 - r) \) \(r - 5 = 25 - 5r\) \(6r = 30\) \(r = 5\) The only solution we get from the equation is r = 5, but we know r = 5 is a restriction on the domain, so there is no solution to this equation. You are correct.

OpenStudy (mathstudent55):

r = 5 is extraneous, and there is no solution to this equation.

OpenStudy (quickstudent):

OK Thanks!